OpenStudy (anonymous):

Factor Completely, 3y^2 +7y + 4

OpenStudy (anonymous):

(3y+4)(y+1)

OpenStudy (anonymous):

(3y + 4) (y + 1)

OpenStudy (anonymous):

do u understand y this is the answe?

OpenStudy (anonymous):

no i dont i cant figure out how to do these

OpenStudy (anonymous):

ok so there is a method

OpenStudy (anonymous):

First see if you can factor but here you can't so we go to the second step

OpenStudy (anonymous):

ax^2+bx+c=y

OpenStudy (anonymous):

Multiply a*c 3*4=12

OpenStudy (anonymous):

so what number when you multiply it equals 12 but when you add it equals b-7?

OpenStudy (anonymous):

So u tell me the answer

OpenStudy (anonymous):

What two numbers when you multiply them equals 12 and when you add them equals 7

OpenStudy (anonymous):

4 and 3

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so now what u wld do is divide the bx into two 7x becomes 3x+4x

OpenStudy (anonymous):

3x+4x=7x so it isn't a big deal to separate them

OpenStudy (anonymous):

So we rewrite the answer question as: 3y^2+3y+4y+4

OpenStudy (anonymous):

Now we separate the equations into two and i cant explain you this steo since i dont get the logic behind it myself

OpenStudy (anonymous):

(3y^2+3y)(4y+4)

OpenStudy (anonymous):

Now factor each bracket 3y(y+1)+4(y+1)

OpenStudy (anonymous):

did you get he steps so far?

OpenStudy (anonymous):

oh u are long gone so i will ust finish off what i was saying

OpenStudy (anonymous):

no im still her just trying to put it all together

OpenStudy (anonymous):

then since 3y and 4 are multiplied by (y+1) we can join them together so (3y+4)(y+1)

OpenStudy (anonymous):

ok basically you gotta memorize the steps and then it becomes second nature